12. If you pruchased 10,000 bushels of soybeans in a futures contract fo $10.00 a bushel, and the price rose to $12.50 prior to maturity, what would your dollar gain or loss be?
a) $2,500 b)$10,000 c)$25,000 d)$15,000
What would your holding period return be assuming you were required to place an initial margin of 10%? a) 250% b) 150% c) 25% d) 350%
Answer #1

Gain = (Selling - buying price)*Number of units

= (12.50-10)*10,000

= $25,000

Hence, the answer is c.

Initial margin = 10*10,000*10% = $10,000

Hence. Holding period return = 25000/10,000

= 250%

i.e. a)